5x+7y=b
mx+y=22
Find the vaule of m and b in the following systems so that the soultion to the system is (-3,4).

Answer by edjones(8007)

(-3,4)
5x+7y=b
mx+y=22 I assume this is the equation.
.
b=5*-3+7*4
=-15+28
b=13
.
-3m+4=22
-3m+4-4=22-4
-3m=18
-3m/-3=18/-3
m=-6
.
Ed